OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12

OpenStudy (zepp):

First of all you need to set the absolute value alone \[|2x+6|=-21\div-3=7\]

OpenStudy (zepp):

Then you'll need to set up two equations Positive ========== \(2x+6=7\) Negative ========== \(2x+6=-7\) Then solve for x! :)

OpenStudy (anonymous):

so its no solution x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No Solution

OpenStudy (zepp):

For positive ======== \(2x+6=7\\2x=7-6=1\\x=\frac{1}{2}\) For negative ======== \(\large 2x+6=-7\\\large 2x=-7-6=-13\\\large x=-\frac{13}{2}\) So none of above, no solutions.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Solve for x: |3x + 2| = 14 how bout htis one